Suppose that a manufacturing firm previously produced at their capacity of 1,000 units. A strong recession is currently causing the firm sell 400 fewer items than before. What is the firm's current capacity utilization?

600

400

40%

60%

1,000
Answer #1

Option 4. 60%

Explanation: Capacity utilization = current production/production capacity * 100 = (1000-400)/1000 * 100 = 600/1000*100 = 0.60*100 = 60%